# Meta-Unlearning on Diffusion Models: Preventing Relearning Unlearned Concepts

### v1

@@ Abstract
With the rapid progress of diffusion-based content generation, significant efforts are being made to unlearn harmful or copyrighted concepts from pretrained diffusion models (DMs) to prevent potential model misuse. However, it is observed that even when DMs are properly unlearned before release, malicious finetuning can compromise this process, causing DMs to relearn the unlearned concepts. This occurs partly because certain benign concepts (e.g., “skin”) retained in DMs are related to the unlearned ones (e.g., “nudity”), facilitating their relearning via finetuning. To address this, we propose meta-unlearning on DMs. Intuitively, a meta-unlearned DM should behave like an unlearned DM when used as is; moreover, if the meta-unlearned DM undergoes malicious finetuning on unlearned concepts, the related benign concepts retained within it will be triggered to self-destruct, hindering the relearning of unlearned concepts. Our meta-unlearning framework is compatible with most existing unlearning methods, requiring only the addition of an easy-to-implement meta objective. We validate our approach through empirical experiments on meta-unlearning concepts from Stable Diffusion models (SD-v1-4 and SDXL), supported by extensive ablation studies. Our code is available at https://github.com/sail-sg/Meta-Unlearning.
@@ 1 Introduction
Diffusion models (DMs) have achieved remarkable success in generative tasks (Ho et al., 2020; Song et al., 2021), leading to the emergence of large-scale models like Stable Diffusion (SD) for text-to-image generation (Rombach et al., 2022). However, training these models often requires vast datasets that may inadvertently contain private or copyrighted content, as well as harmful concepts that are not safe for work (NSFW) (Schramowski et al., 2023). These challenges have sparked interest in machine unlearning algorithms for DMs (Gandikota et al., 2023; 2024; Kumari et al., 2023; Kim et al., 2023), which modify pretrained models to forget specific inappropriate data (forget set) while retaining performance on the remaining benign data (retain set).
While unlearning methods designed for DMs have shown promising results, recent studies reveal that unlearned models may be maliciously induced to relearn the unlearned concepts during finetuning, even when the finetuning is performed on unrelated benign data (Qi et al., 2023; Tamirisa et al., 2024; Patil et al., 2024; Shumailov et al., 2024). Although these studies focus primarily on language models, we observe similar phenomena on DMs as shown in Fig. 3. This partly occurs because certain benign concepts (e.g., “skin” in the retain set) related to unlearned ones (e.g., “nudity” in the forget set) are still retained in DMs, easing their relearning during finetuning.
To tackle this challenge, we draw inspiration from meta-learning (Finn et al., 2017) and propose the meta-unlearning framework. This framework comprises two key components: (1) a standard unlearning objective to ensure the model effectively forgets specified data before public release, while preserving performance on benign data; and (2) a meta objective designed to slow down the relearning process if the model is maliciously finetuned on the forget set. Additionally, it induces the benign knowledge related to the forget set to self-destruct, as illustrated in Fig. 1.
Our meta-unlearning framework is compatible with most existing unlearning methods for DMs, requiring only the addition of a simple-to-implement meta objective, as outlined in Algorithm 1. This meta objective can be efficiently optimized by automatic differentiation (Paszke et al., 2019). We conduct extensive experiments on SD models (SD-v1-4 and SDXL) to validate the effectiveness of various instantiations of our meta-unlearning approach.
@@ 2 Related work
Recent studies have shown that DMs can be misused to generate unsafe content, such as images depicting sexual acts, harassment, or illegal activities (Schramowski et al., 2023; Gao et al., 2023; Rando et al., 2022). To mitigate this issue, early-stage DMs are equipped with NSFW filters designed to block the generation of inappropriate images (Rando et al., 2022). However, this approach does not prevent the model from generating harmful imagery at its core, and these filters can be easily bypassed, exposing significant security vulnerabilities (Birhane et al., 2021; Rombach et al., 2022). As a result, machine unlearning methods have been proposed for DMs.
2.1 Machine unlearning on DMs
@@
Several methods have been proposed to unlearn or erase harmful, private, or copyrighted concepts from DMs (Zhang et al., 2024a; c; Park et al., 2024; Huang et al., 2023; Wu et al., 2024; Pham et al., 2024b; Lu et al., 2024). For instances, ESD (Gandikota et al., 2023) leverages negative guidance to finetune the U-Net, removing the specified style or concept. Concept ablation (Kumari et al., 2023) works by making the distribution of the target concept similar to that of an anchor concept. However, these methods are vulnerable to adversarial attacks. To this end, several adversarial-resistant unlearning methods have been proposed (Li et al., 2024b; Yang et al., 2024; Kim et al., 2024; Huang et al., 2024b). AdvUnlearn (Zhang et al., 2024c) enhances the robustness of concept erasure by incorporating adversarial training principles, while RECE (Gong et al., 2024) derives new target embeddings for inappropriate content and iteratively aligns them with harmless concepts in cross-attention layers. Despite these advancements, the models unlearned by these methods can still be maliciously finetuned to relearn unlearned concepts, as observed in our experiments.
2.2 Machine unlearning on language models
While this paper primarily focuses on unlearning DMs, there have been a lot of efforts devoted to unlearning language models (Yao et al., 2023; Maini et al., 2024; Wang et al., 2024b; Li et al., 2024a; Yao et al., 2024; Gu et al., 2024; Zhang et al., 2024b; Jia et al., 2024; Tian et al., 2024; Tang et al., 2024; Tamirisa et al., 2024). These methods typically finetune the model on a forget set. In addition, there are also other tuning-free unlearning techniques, including contrastive decoding (Huang et al., 2024a; Wang et al., 2024a; Ji et al., 2024; Dong et al., 2024), task vectors (Dou et al., 2024; Liu et al., 2024b), in-context learning (Pawelczyk et al., 2023; Muresanu et al., 2024; Thaker et al., 2024), and input processing and detection (Bhaila et al., 2024; Gao et al., 2024; Liu et al., 2024a).
2.3 Unlearned models can be attacked
@@
Recent studies have also demonstrated that unlearned models are vulnerable to generating previously unlearned concepts through adversarial attacks (Zhang et al., 2023; Tsai et al., 2023; Pham et al., 2024a; Ma et al., 2024) and malicious finetuning (Tamirisa et al., 2024; Shumailov et al., 2024; Łucki et al., 2024; Qi et al., 2023). For instance, UnlearnDiffAtk (Zhang et al., 2023) introduces an evaluation framework that uses adversarial attacks to generate adversarial prompts by exploiting the inherent classification capabilities of DMs. In the domain of language models, several works have revealed that finetuning can recover unlearned concepts. For example, Qi et al. (2023) demonstrate that safety alignment and/or unlearning in language models can be undermined through finetuning with a small set of adversarially crafted training examples. Additionally, Tamirisa et al. (2024) show that refusal mechanisms and unlearning safeguards can be bypassed with minimal iterations of finetuning, while Łucki et al. (2024) recover most supposedly unlearned capabilities.
@@ 3 Preliminaries
This section provides a brief review of diffusion models (DMs) (Ho et al., 2020; Song et al., 2021) and commonly used machine unlearning methods in the DM literature.
@@ 3.1 Diffusion Models
Our research focuses on discrete-time DMs, especially latent diffusion models (LDMs) that serve as the cornerstone of Stable Diffusion (Rombach et al., 2022). We consider random variables $\bm{x}\in\mathcal{X}$ and $\bm{c}\in\mathcal{C}$, where $\bm{x}$ denotes the latent feature and $\bm{c}$ the conditional context, e.g., text prompts. Let $q(\bm{x},\bm{c})$ denote the data distribution. Consider a forward diffusion process over time interval $[0,T]$ with $T\in\mathbb{N}^{+}$. The Markov transition probability from $\bm{x}_{t-1}$ to $\bm{x}_{t}$ is $q(\bm{x}_{t}|\bm{x}_{t-1})\triangleq\mathcal{N}(\bm{x}_{t}|\sqrt{1-\beta_{t}}\bm{x}_{t-1},\beta_{t}\mathbf{I})$, where $\bm{x}_{0}=\bm{x}$ and $\beta_{1},\cdots,\beta_{T}$ correspond to a variance schedule. Note that we can sample $\bm{x}_{t}$ at an arbitrary timestep $t$ directly from $\bm{x}$, since there is $q(\bm{x}_{t}|\bm{x})=\mathcal{N}(\bm{x}_{t}|\sqrt{\overline{\alpha}_{t}}\bm{x},(1-\overline{\alpha}_{t})\mathbf{I})$, where $\alpha_{t}\triangleq 1-\beta_{t}$ and $\overline{\alpha}_{t}\triangleq\prod_{i=1}^{t}\alpha_{i}$.
Sohl-Dickstein et al. (2015) show that when $\beta_{t}$ are small, the reverse diffusion process can also be modeled by Gaussian conditionals. Specifically, the reverse transition probability from $\bm{x}_{t}$ to $\bm{x}_{t-1}$ is written as $p_{\theta}(\bm{x}_{t-1}|\bm{x}_{t},\bm{c})=\mathcal{N}(\bm{x}_{t-1}|\bm{\mu}_{\theta}(\bm{x}_{t},\bm{c}),\sigma_{t}^{2}\mathbf{I})$, where $\theta\in\mathbb{R}^{d}$ is the model parameters and $\sigma_{t}$ are time dependent constants. Instead of directly modeling the data prediction $\bm{\mu}_{\theta}$, we choose to model the noise prediction $\bm{\epsilon}_{\theta}$ based on the parameterization $\bm{\mu}_{\theta}(\bm{x}_{t},\bm{c})=\frac{1}{\sqrt{\alpha_{t}}}\left(\bm{x}_{t}-\frac{\beta_{t}}{\sqrt{1-\overline{\alpha}_{t}}}\bm{\epsilon}_{\theta}(\bm{x}_{t},\bm{c})\right)$. The training objective of $\bm{\epsilon}_{\theta}(\bm{x}_{t},\bm{c})$ can be derived from optimizing the (weighted) variational bound of negative log-likelihood, formulated as follows:
$\min_{\theta}\mathcal{L}_{\textrm{DM}}(\theta;\mathcal{D}_{\textrm{train}})=\mathbb{E}_{(\bm{x},\bm{c})\sim\mathcal{D}_{\textrm{train}},\bm{\epsilon},t}\left[\left\|\bm{\epsilon}-\bm{\epsilon}_{\theta}(\bm{x}_{t},\bm{c})\right\|_{2}^{2}\right]\textrm{,}$ (1)
where $\bm{x}_{t}=\sqrt{\overline{\alpha}_{t}}\bm{x}+\sqrt{1-\overline{\alpha}_{t}}\bm{\epsilon}$, the data pairs $(\bm{x},\bm{c})$ are sampled from the training dataset $\mathcal{D}_{\textrm{train}}$, $\bm{\epsilon}\sim\mathcal{N}(\bm{\epsilon}|\mathbf{0},\mathbf{I})$ is a standard Gaussian noise, and $t\sim\mathcal{U}([1,T])$ follows the uniform distribution.
@@ 3.2 Machine Unlearning for DMs
DMs, despite their high capability, may generate unsafe content or disclose sensitive information that is not safe for work (NSFW) (Schramowski et al., 2023). Several recent studies have investigated concept erasing or machine unlearning for DMs to address safety, privacy, and copyright concerns (Kumari et al., 2023; Zhang et al., 2024c; Heng & Soh, 2024). Let $\bm{\epsilon}_{\theta^{*}}$ denotes the DM pretrained on the dataset $\mathcal{D}_{\textrm{train}}$, where $\theta^{*}=\arg\min_{\theta}\mathcal{L}_{\textrm{DM}}(\theta;\mathcal{D}_{\textrm{train}})$. The goal of machine unlearning is to unlearn a forget set $\mathcal{D}_{\text{forget}}\subset\mathcal{D}_{\textrm{train}}$ from $\bm{\epsilon}_{\theta^{*}}$, while preserving performance on the retain set $\mathcal{D}_{\text{retain}}=\mathcal{D}_{\text{train}}\backslash\mathcal{D}_{\text{forget}}$. We describe four unlearning methods for DMs that we use as baselines:
@@
  • Erased Stable Diffusion (ESD) (Gandikota et al., 2023) intervenes pretrained DMs by steering generation away from the concept intended to be forgotten. Ideally, the unlearned DM is expected to predict $\widetilde{\bm{\epsilon}}_{\theta^{*}}(\bm{x}_{t},\bm{c})=\bm{\epsilon}_{\theta^{*}}(\bm{x}_{t},\emptyset)-\eta\left[\bm{\epsilon}_{\theta^{*}}(\bm{x}_{t},\bm{c})-\bm{\epsilon}_{\theta^{*}}(\bm{x}_{t},\emptyset)\right]$ when fed in $(\bm{x},\bm{c})\sim\mathcal{D}_{\textrm{forget}}$, where $\eta>0$ is a hyperparameter and $\emptyset$ indicates unconditional context. The unlearning objective of ESD is to optimize
@@
    $\min_{\theta}\mathcal{L}_{\textrm{ESD}}(\theta;\mathcal{D}_{\textrm{forget}})=\mathbb{E}_{(\bm{x},\bm{c})\sim\mathcal{D}_{\textrm{forget}},\bm{\epsilon},t}\left[\left\|\bm{\epsilon}_{\theta}(\bm{x}_{t},\bm{c})-\widetilde{\bm{\epsilon}}_{\theta^{*}}(\bm{x}_{t},\bm{c})\right\|_{2}^{2}\right]\textrm{,}$ (2)
@@
    where $\theta$ is initialized from the frozen $\theta^{*}$. Gandikota et al. (2023) use ESD-x-$\eta$ to indicate only cross-attention parameters are finetuned with hyperparameter $\eta$; likewise, ESD-u-$\eta$ indicates only non-cross-attention parameters are finetuned, and ESD-f-$\eta$ indicates full finetuning.
@@
  • Safe self-distillation diffusion (SDD) (Kim et al., 2023) is a self-distillation paradigm to erase concepts from DMs. The unlearning objective of SDD is to optimize
@@
    $\min_{\theta}\mathcal{L}_{\textrm{SDD}}(\theta;\mathcal{D}_{\textrm{forget}})=\mathbb{E}_{(\bm{x},\bm{c})\sim\mathcal{D}_{\textrm{forget}},\bm{\epsilon},t}\left[\left\|\bm{\epsilon}_{\theta}(\bm{x}_{t},\bm{c})-\texttt{sg}\left(\bm{\epsilon}_{\theta}(\bm{x}_{t},\emptyset)\right)\right\|_{2}^{2}\right]\textrm{,}$ (3)
@@
    where sg is the stop-gradient operation and $\theta$ is initialized from the frozen $\theta^{*}$. To mitigate catastrophic forgetting, SDD employs an exponential moving average (EMA) teacher. Note that in the original implementations of both ESD and SDD, there are only text prompts $\bm{c}$ in the forget set, while the noisy latents $\bm{x}_{t}$ are generated by the frozen DM $\bm{\epsilon}_{\theta^{*}}$.
@@
  • Unified concept editing (UCE) (Gandikota et al., 2024) edits the pretrained DMs via a closed-form solution without finetuning. Let $W^{*}$ be the attention matrices of $\theta^{*}$ (Key/Value matrices), $\mathcal{T}$ be the text embedding mapping in $\bm{\epsilon}_{\theta^{*}}$, then the unlearning objective of UCE is to optimize
@@
    $\!\!\!\!\min_{W}\mathbb{E}_{\bm{c}_{f},\bm{c}_{r}}\left[\left\|W\mathcal{T}(\bm{c}_{f})-W^{*}\mathcal{T}(\emptyset)\right\|_{2}^{2}+\lambda_{1}\left\|W\mathcal{T}(\bm{c}_{r})-W^{*}\mathcal{T}(\bm{c}_{r})\right\|_{2}^{2}+\lambda_{2}\left\|W-W^{*}\right\|_{2}^{2}\right]\!\textrm{,}\!\!$ (4)
@@
    where $\bm{c}_{f}\sim\mathcal{D}_{\textrm{forget}}$, $\bm{c}_{r}\sim\mathcal{D}_{\textrm{retain}}$, and $\lambda_{1}$, $\lambda_{2}$ are hyperparameters. Gandikota et al. (2024) prove that the above minimization problem has closed-form solution $W_{\textrm{UCE}}$.
@@
  • Reliable and efficient concept erasure (RECE) (Gong et al., 2024) first performs UCE, after which iteratively creates new erasing embeddings and obtains updated attention matrices. Specifically, let $\widetilde{W}\leftarrow W_{\textrm{UCE}}$, and use subscripts $i$ to denote the $i$-th attention matrix in the model; then RECE iteratively constructs $\bm{c}^{\prime}$ by optimizing
    $\min_{\bm{c}^{\prime}}\sum_{i}\left\|\widetilde{W}_{i}\mathcal{T}(\bm{c}^{\prime})-W_{i}^{*}\mathcal{T}(\bm{c}_{f})\right\|_{2}^{2}+\lambda\left\|\mathcal{T}(\bm{c}^{\prime})\right\|_{2}^{2}\textrm{,}$ (5)

### v2

@@ Abstract
With the rapid progress of diffusion models (DMs), significant efforts are being made to unlearn harmful or copyrighted concepts from pretrained DMs to prevent potential model misuse. However, it is observed that even when DMs are properly unlearned before release, malicious finetuning can compromise this process, causing DMs to relearn the unlearned concepts. This occurs partly because certain benign concepts (e.g., “skin”) retained in DMs are related to the unlearned ones (e.g., “nudity”), facilitating their relearning via finetuning. To address this, we propose meta-unlearning on DMs. Intuitively, a meta-unlearned DM should behave like an unlearned DM when used as is; moreover, if the meta-unlearned DM undergoes malicious finetuning on unlearned concepts, the related benign concepts retained within it will be triggered to self-destruct, hindering the relearning of unlearned concepts. Our meta-unlearning framework is compatible with most existing unlearning methods, requiring only the addition of an easy-to-implement meta objective. We validate our approach through empirical experiments on meta-unlearning concepts from Stable Diffusion models (SD-v1-4 and SDXL), supported by extensive ablation studies.
@@ 1 Introduction
Diffusion models (DMs) have achieved remarkable success in generative tasks [17, 58], leading to the emergence of large-scale models like Stable Diffusion (SD) for text-to-image generation [51]. However, training these models often requires vast datasets that may inadvertently contain private or copyrighted content, as well as harmful concepts that are not safe for work (NSFW) [54]. These challenges have sparked interest in machine unlearning algorithms for DMs [6, 7, 25, 24], which modify pretrained models to forget specific inappropriate data (forget set) while retaining performance on the remaining benign data (retain set).
While unlearning methods designed for DMs are promising, recent studies reveal that unlearned models may be maliciously finetuned to relearn the unlearned concepts, even when the finetuning is performed on unrelated benign data [48, 59, 42, 55]. Although these studies focus primarily on language models, we observe similar phenomena on DMs as shown in Fig. 4. This partly occurs because certain benign concepts (e.g., “skin” in the retain set) related to unlearned ones (e.g., “nudity” in the forget set) are still retained in DMs, easing their relearning during finetuning.
To tackle this challenge, we draw inspiration from meta-learning [5] and propose the meta-unlearning framework. This framework comprises two key components: (1) a standard unlearning objective to ensure the model effectively forgets specified data before public release, while preserving performance on benign data; and (2) a meta objective designed to slow down the relearning process if the model is maliciously finetuned on the forget set. Additionally, it induces the benign knowledge related to the forget set to self-destruct, as illustrated in Fig. 1.
Our meta-unlearning framework is compatible with most existing unlearning methods for DMs, requiring only the addition of a simple-to-implement meta objective, as outlined in Algorithm 1. This meta objective can be efficiently optimized by automatic differentiation [41]. We conduct extensive experiments on SD models (SD-v1-4 and SDXL) to validate the effectiveness of various instantiations of our meta-unlearning approach.
@@ 2 Related work
Recent studies have shown that DMs can be misused to generate unsafe content, such as images depicting sexual acts, harassment, or illegal activities [54, 9, 50]. To mitigate this issue, early-stage DMs are equipped with NSFW filters designed to block the generation of inappropriate images [50]. However, this approach does not prevent the model from generating harmful imagery at its core, and these filters can be bypassed, exposing security vulnerabilities [2, 51].
@@
Machine unlearning on DMs. Several methods have been proposed to unlearn or erase harmful, private, or copyrighted concepts from DMs [71, 74, 40, 18, 66, 46, 33]. For instances, ESD [6] leverages negative guidance to finetune the U-Net, removing the specified style or concept. Concept ablation [25] works by making the distribution of the target concept similar to that of an anchor concept. However, these methods are vulnerable to adversarial attacks. To this end, several adversarial-resistant unlearning methods have been proposed [29, 67, 23, 20]. AdvUnlearn [74] enhances the robustness of concept erasure by incorporating adversarial training principles, while RECE [10] derives new target embeddings for inappropriate content and iteratively aligns them with harmless concepts in cross-attention layers. Nevertheless, models unlearned by these methods can still be maliciously finetuned to relearn unlearned concepts, as observed in our experiments.
@@
Unlearned models can be attacked. Recent studies have shown that unlearned models are vulnerable to generating previously unlearned concepts through adversarial attacks [73, 63, 45, 34, 44, 12, 70] and malicious finetuning [59, 55, 75, 48]. UnlearnDiffAtk [73] introduces an evaluation framework that uses adversarial attacks to generate adversarial prompts by exploiting the inherent classification capabilities of DMs. In the domain of language models, several works have revealed that finetuning can recover unlearned concepts. For example, [48] demonstrates that safety alignment and/or unlearning in language models can be undermined through finetuning with a small set of adversarially crafted training examples. Additionally, [59] show that refusal mechanisms and unlearning safeguards can be bypassed with minimal iterations of finetuning, while [75] recover most supposedly unlearned capabilities.
@@ 3 Preliminaries
This section provides a brief review of diffusion models (DMs) [17, 58] and commonly used machine unlearning methods in the DM literature.
@@ 3.1 Diffusion Models
We focus on discrete-time DMs that serve as the cornerstone of Stable Diffusion [51]. We consider random variables $\bm{x}\in\mathcal{X}$ and $\bm{c}\in\mathcal{C}$, where $\bm{x}$ denotes the latent feature and $\bm{c}$ the conditional context, e.g., text prompts. Let $q(\bm{x},\bm{c})$ denote the data distribution. Consider a forward diffusion process over time interval $[0,T]$ with $T\in\mathbb{N}^{+}$. The Markov transition probability from $\bm{x}_{t-1}$ to $\bm{x}_{t}$ is $q(\bm{x}_{t}|\bm{x}_{t-1})\triangleq\mathcal{N}(\bm{x}_{t}|\sqrt{1-\beta_{t}}\bm{x}_{t-1},\beta_{t}\mathbf{I})$, where $\bm{x}_{0}=\bm{x}$ and $\beta_{1},\cdots,\beta_{T}$ correspond to a variance schedule. Note that we can sample $\bm{x}_{t}$ at an arbitrary timestep $t$ directly from $\bm{x}$, since there is $q(\bm{x}_{t}|\bm{x})=\mathcal{N}(\bm{x}_{t}|\sqrt{\overline{\alpha}_{t}}\bm{x},(1-\overline{\alpha}_{t})\mathbf{I})$, where $\alpha_{t}\triangleq 1-\beta_{t}$ and $\overline{\alpha}_{t}\triangleq\prod_{i=1}^{t}\alpha_{i}$.
Sohl-Dickstein et al. [57] show that when $\beta_{t}$ are small, the reverse diffusion process can also be modeled by Gaussian conditionals. Specifically, the reverse transition probability from $\bm{x}_{t}$ to $\bm{x}_{t-1}$ is written as $p_{\theta}(\bm{x}_{t-1}|\bm{x}_{t},\bm{c})=\mathcal{N}(\bm{x}_{t-1}|\bm{\mu}_{\theta}(\bm{x}_{t},\bm{c}),\sigma_{t}^{2}\mathbf{I})$, where $\theta\in\mathbb{R}^{d}$ is the model parameters and $\sigma_{t}$ are time dependent constants. Instead of directly modeling the data prediction $\bm{\mu}_{\theta}$, we choose to model the noise prediction $\bm{\epsilon}_{\theta}$ based on the parameterization $\bm{\mu}_{\theta}(\bm{x}_{t},\bm{c})=\frac{1}{\sqrt{\alpha_{t}}}\left(\bm{x}_{t}-\frac{\beta_{t}}{\sqrt{1-\overline{\alpha}_{t}}}\bm{\epsilon}_{\theta}(\bm{x}_{t},\bm{c})\right)$. The training objective of $\bm{\epsilon}_{\theta}(\bm{x}_{t},\bm{c})$ can be derived from optimizing the (weighted) variational bound of negative log-likelihood, formulated as follows:
$\min_{\theta}\mathcal{L}_{\textrm{DM}}(\theta;\mathcal{D}_{\textrm{train}})=\mathbb{E}_{(\bm{x},\bm{c})\sim\mathcal{D}_{\textrm{train}},\bm{\epsilon},t}\left[\left\|\bm{\epsilon}-\bm{\epsilon}_{\theta}(\bm{x}_{t},\bm{c})\right\|_{2}^{2}\right]\textrm{,}$
where $\bm{x}_{t}=\sqrt{\overline{\alpha}_{t}}\bm{x}+\sqrt{1-\overline{\alpha}_{t}}\bm{\epsilon}$, the pairs $(\bm{x},\bm{c})$ are sampled from the training set $\mathcal{D}_{\textrm{train}}$, $\bm{\epsilon}\sim\mathcal{N}(\bm{\epsilon}|\mathbf{0},\mathbf{I})$ is a standard Gaussian noise, and $t\sim\mathcal{U}([1,T])$ is a uniform distribution.
@@ 3.2 Machine Unlearning for DMs
DMs, despite their high capability, may generate unsafe content or disclose sensitive information that is not safe for work (NSFW) [54]. Several recent studies have investigated concept erasing or machine unlearning for DMs to address safety, privacy, and copyright concerns [25, 74, 14]. Let $\bm{\epsilon}_{\theta^{*}}$ denotes the DM pretrained on the dataset $\mathcal{D}_{\textrm{train}}$, where $\theta^{*}=\arg\min_{\theta}\mathcal{L}_{\textrm{DM}}(\theta;\mathcal{D}_{\textrm{train}})$. The goal of machine unlearning is to unlearn a forget set $\mathcal{D}_{\text{forget}}\subset\mathcal{D}_{\textrm{train}}$ from $\bm{\epsilon}_{\theta^{*}}$, while preserving performance on the retain set $\mathcal{D}_{\text{retain}}=\mathcal{D}_{\text{train}}\backslash\mathcal{D}_{\text{forget}}$. Below we describe four unlearning methods for DMs that we use as baselines.
@@
Erased Stable Diffusion (ESD) [6] intervenes pretrained DMs by steering generation away from the concept intended to be forgotten. Ideally, the unlearned DM is expected to predict $\widetilde{\bm{\epsilon}}_{\theta^{*}}(\bm{x}_{t},\bm{c})=\bm{\epsilon}_{\theta^{*}}(\bm{x}_{t},\emptyset)-\eta\left[\bm{\epsilon}_{\theta^{*}}(\bm{x}_{t},\bm{c})-\bm{\epsilon}_{\theta^{*}}(\bm{x}_{t},\emptyset)\right]$ when fed in $(\bm{x},\bm{c})\sim\mathcal{D}_{\textrm{forget}}$, where $\eta>0$ is a hyperparameter and $\emptyset$ indicates unconditional context. The unlearning objective of ESD is
@@
$\begin{split}&\min_{\theta}\mathcal{L}_{\textrm{ESD}}(\theta;\mathcal{D}_{\textrm{forget}})\\ =&\mathbb{E}_{(\bm{x},\bm{c})\sim\mathcal{D}_{\textrm{forget}},\bm{\epsilon},t}\left[\left\|\bm{\epsilon}_{\theta}(\bm{x}_{t},\bm{c})-\widetilde{\bm{\epsilon}}_{\theta^{*}}(\bm{x}_{t},\bm{c})\right\|_{2}^{2}\right]\textrm{,}\end{split}$
@@
where $\theta$ is initialized from the frozen $\theta^{*}$. Gandikota et al. [6] use ESD-x-$\eta$ to indicate only cross-attention parameters are finetuned with hyperparameter $\eta$; likewise, ESD-u-$\eta$ indicates only non-cross-attention parameters are finetuned, and ESD-f-$\eta$ indicates full finetuning.
@@
Safe self-distillation diffusion (SDD) [24] is a self-distillation paradigm to erase concepts from DMs. The unlearning objective of SDD is to optimize
@@
$\begin{split}&\min_{\theta}\mathcal{L}_{\textrm{SDD}}(\theta;\mathcal{D}_{\textrm{forget}})\\ =&\mathbb{E}_{(\bm{x},\bm{c})\sim\mathcal{D}_{\textrm{forget}},\bm{\epsilon},t}\left[\left\|\bm{\epsilon}_{\theta}(\bm{x}_{t},\bm{c})-\texttt{sg}\left(\bm{\epsilon}_{\theta}(\bm{x}_{t},\emptyset)\right)\right\|_{2}^{2}\right]\textrm{,}\end{split}$
@@
where sg is the stop-gradient operation and $\theta$ is initialized from the frozen $\theta^{*}$. To mitigate catastrophic forgetting, SDD employs an EMA teacher. Note that in the original implementations of both ESD and SDD, there are only text prompts $\bm{c}$ in the forget set, while the noisy latents $\bm{x}_{t}$ are generated by the frozen DM $\bm{\epsilon}_{\theta^{*}}$.
@@
Unified concept editing (UCE) [7] edits the pretrained DMs via a closed-form solution without finetuning. Let $W^{*}$ be the attention matrices of $\theta^{*}$ (Key/Value matrices), $\mathcal{T}$ be the text embedding mapping in $\bm{\epsilon}_{\theta^{*}}$, then the unlearning objective of UCE is to optimize
@@
$\begin{split}\min_{W}&\mathbb{E}_{\bm{c}_{f},\bm{c}_{r}}\Big{[}\left\|W\mathcal{T}(\bm{c}_{f})-W^{*}\mathcal{T}(\emptyset)\right\|_{2}^{2}+\\ &\lambda_{1}\left\|W\mathcal{T}(\bm{c}_{r})-W^{*}\mathcal{T}(\bm{c}_{r})\right\|_{2}^{2}+\lambda_{2}\left\|W-W^{*}\right\|_{2}^{2}\Big{]}\textrm{,}\end{split}$
@@
where $\bm{c}_{f}\sim\mathcal{D}_{\textrm{forget}}$, $\bm{c}_{r}\sim\mathcal{D}_{\textrm{retain}}$, and $\lambda_{1}$, $\lambda_{2}$ are hyperparameters. Gandikota et al. [7] prove that the above minimization problem has closed-form solution $W_{\textrm{UCE}}$.
@@
Reliable and efficient concept erasure (RECE) [10] first performs UCE, after which iteratively creates new erasing embeddings and obtains updated attention matrices. Specifically, let $\widetilde{W}\leftarrow W_{\textrm{UCE}}$, and use subscripts $i$ to denote the $i$-th attention matrix in the model; then RECE iteratively constructs $\bm{c}^{\prime}$ by optimizing
$\min_{\bm{c}^{\prime}}\sum_{i}\left\|\widetilde{W}_{i}\mathcal{T}(\bm{c}^{\prime})-W_{i}^{*}\mathcal{T}(\bm{c}_{f})\right\|_{2}^{2}+\lambda\left\|\mathcal{T}(\bm{c}^{\prime})\right\|_{2}^{2}\textrm{,}$
@@
Algorithm 1 The general framework of meta-unlearning
1:Pretrained $\theta^{*}$, forget set $\mathcal{D}_{\text{forget}}$, retain set $\mathcal{D}_{\text{retain}}$
3:Outer (steps NN, learning rate $\omega$), inner (steps $M$, learning rate $\tau$), scale factors $\gamma_{1},\gamma_{2}$
4:# If $\mathcal{L}_{\textrm{unlearn}}$ is ESD/SDD that needs optimization
[…]
6:# If $\mathcal{L}_{\textrm{unlearn}}$ is UCE/RECE that has closed-form solution
[…]
8:for $n=1$ to NN do
9:  Sample a finetuning set $\mathcal{D}_{\text{FT}}\subset\mathcal{D}_{\text{forget}}$
10:  Initialize $\bm{g}=\bm{0}$ and $\theta^{\textrm{FT}}=\theta_{n-1}$
11:   # If $\mathcal{L}_{\textrm{unlearn}}$ is ESD/SDD
12:  $\bm{g}\leftarrow\bm{g}+\gamma_{1}\cdot\nabla_{\theta_{n-1}}\mathcal{L}_{\text{unlearn}}(\theta_{n-1};\mathcal{D}_{\text{forget}},\mathcal{D}_{\text{retain}})$
13:  for $m=1$ to $M$ do
14:   $\theta^{\textrm{FT}}\leftarrow\theta^{\textrm{FT}}-\tau\cdot\nabla_{\theta^{\textrm{FT}}}\mathcal{L}_{\textrm{FT}}\left(\theta;\mathcal{D}_{\textrm{FT}}\right)$
15:  end for
16:   # Meta objective
17:  $\bm{g}\leftarrow\bm{g}+\gamma_{2}\cdot\nabla_{\theta_{n-1}}\mathcal{L}_{\text{meta}}(\theta^{\textrm{FT}};\mathcal{D}_{\text{FT}},\mathcal{D}_{\text{retain}})$
18:  $\theta_{n}\leftarrow\theta_{n-1}-\omega\cdot\bm{g}$
19:end for
20:return $\theta_{N}$